Convert 0.4 radians into degrees.
Found 2 solutions by Alan3354, oberobic:
Answer by Alan3354(69443)

Convert 0.4 radians into degrees.
---------
0.4 rad = 0.4*180/pi degs
=~ 22.9 degs

Answer by oberobic(2304)

The simple rule is that there are 2*pi radians in 360 degrees.
.
Solve using proportions.
.
.4 / 2*pi = x / 360
.
Cross multiply.
.
.4*360 = 2*pi*x
.
Divide both sides by the constant 2*pi.
.
x = (.4*360) / 2*pi = 144/(2*pi) = 72/pi = 22.91831180523293
.
.4 radians = 22.92 degrees
.
To check the answer, you can type "convert .4 radians to degrees" into the Google search engine box.
.
Done.